OpenStudy (anonymous):

A bag contains green, blue and yellow balls. The ratio of green to blue balls is 2:7. The ratio of green to yellow balls is 3:5. Quantity A Quantity B Number of blue balls Number of yellow balls

OpenStudy (anonymous):

I FOUND THE SOLUTION ONLINE BUT CAN'T UNDERSTAND HOW THEY CAME UP WITH B/G Solution Let g, b and y be the numbers of green, blue and yellow balls respectively. Using the given ratios, we can write the following fractions g / b = 2 / 7 and g / y = 3 / 5 The first fraction gives b / g = 7 / 2 We now evaluate the product of fractions g / y and b / g as follows (g / y) * (b / g) = (3 / 5) * (7 / 2) Note that (g / y) * (b / g) simplifies to b / y, hence b / y

OpenStudy (anonymous):

??

OpenStudy (mathmate):

 I don't know the total number of balls, so cannot solve in numbers. The above diagram shows the ratio among the three colours.